OpenStudy (anonymous):

what is the square root of 200

OpenStudy (anonymous):

14

OpenStudy (bibby):

you can write 200 as 4*50 and further simplify that as 4*25*2 from there use the property that \(\sqrt{a*b*c}=\sqrt a*\sqrt b*\sqrt c\)

OpenStudy (anonymous):

so wat is the answer

OpenStudy (hungergames13):

14

OpenStudy (bibby):

simplify \(\sqrt{100} * \sqrt{2} \) or \(\sqrt4*\sqrt{25}*\sqrt2\)

OpenStudy (bibby):

\(10\sqrt2\)